# Rough Sets for Explainability of Spectral Graph Clustering

### v1

@@ 1 Introduction
Explainability turns out to be crucial for the application of Machine Learning techniques, in particular clustering. From this perspective, Graph Spectral Clustering (GSC for short) methods have some serious drawbacks as they operate in a space hard to understand by humans, that is, in the Laplacian spectral domain, and this is why they are considered opaque models [^1] See also, for example [Sabbatini:2022], and https://crunchingthedata.com/when-to-use-spectral-clustering/. . Model-unaware methods of textual clustering are known and could be considered an option, but they only describe groups and do not explain why documents are included. Recently explanatory clustering methods for Combinatorial and Normalized Laplacian-based GSC were proposed for the textual domain, as presented in [Plosone2025]. Yet, two problems remain open: (1) in real-world settings, there exist objects that do not fit any cluster and are forced into some clusters just distorting them, and (2) clustering methods do not achieve the real optimal value of the quality function so that some cluster elements may also not fit the concepts and can distort them. In this paper, we want to contribute to solving both problems by applying concepts from rough set theory, where entities, e.g. documents, are divided into core and boundary objects.
@@
Thus, the approach introduced in the context of rough sets by [Paw82] may prove fruitful here as it allows one to distinguish between cluster core elements and boundary elements. This distinction can, for example, result from repeated runs of a particular probabilistic clustering algorithm. Elements that belong to the same cluster on each run are (with high probability) the core of a cluster, while the others can be considered to belong to the boundaries of different clusters. Another situation, though bearing some similarity, is when we apply different algorithms optimizing the same criterion, but each in an approximate way. The elements fitting together may be considered the core, with a high probability of fitting the (approximated) criterion, while the other will constitute boundaries of appropriate clusters. However, another approach can be a robustness study, checking the impact on the optimality criterion of randomly shuffling parts of the established clusters.
@@ 2 Brief introduction to Graph Spectral Analysis
There are several graph clustering methods; see, for example, [Schaeffer], aimed at discovering clusters that meet related but slightly different criteria. The best known are the RCut criterion and the NCut criterion.
Assume that the matrix $S$, called the similarity matrix, represents similarities between pairs of $n$ objects with entries ranging from 0 to 1. $S$ is assumed to have a diagonal equal to zero, as self-loops are not allowed (see, e.g. [TU:2022:3673]).
@@
where $n_{j}=|C_{j}|$. A detailed discussion leading to the above formulation can be found, for example, in [Plosone2025].
@@
where $\mathcal{V}_{j}=\sum_{i\in C_{j}}d_{ii}$ denotes the volume of the node $j$, and $d_{ii}=\sum_{j=1}^{n}s_{ij}$ stands for the degree of the node $i$. Thus, we try to minimize the total weights of links joining nodes belonging to opposite clusters. In the first case, this sum is divided by the cardinalities of the respective clusters, while in the second case - by the volumes of these clusters. Again, see [Plosone2025] for a more in-depth discussion.
@@
where $D$ is the degree matrix, i.e., a diagonal matrix whose elements are the degrees $d_{jj}$ of all nodes. A normalized Laplacian [^2] Other Laplacians are also used (see, [STWMAKSpringer:2018]), e.g. the random walk Laplacian $\mathbb{L}$ of a graph, defined as $\mathbb{L}=LD^{-1}=I-SD^{-1}$ (5) $\mathcal{L}$ of the graph represented by $S$ is defined as
@@
Based on the Rayleigh-Ritz theorem [Gol96], it is concluded that the sought pseudo-indicator vectors are the eigenvectors of the respective Laplacian, related to the lowest eigenvalues. If we cluster into $k$ clusters, we take the lowest $k$ eigenvectors (except the last, which is certain to be equal to zero). These pseudo-indicator vectors span the embedding space for the objects which are then clustered in this embedding using, e.g. $k$-means algorithm. For an explanation and further details, see, e.g. [vonLuxburg:2007] or [STWMAKSpringer:2018]. Afterward, note that $L$-based clustering means GSC based on combinatorial Laplacian $L$, and $N$-based clustering means GSC based on normalized Laplacian $\mathfrak{L}$. Note that by projecting the original data onto the space spanned by $k$ eigenvectors of the corresponding Laplacian, we discover an interesting feature of such a procedure: it naturally combines clustering with (nonlinear) dimensionality reduction.
@@ 2.1 $K$-embedding
The $K$-embedding [Plosone2025] aims at placing objects $i,\ell$ at distances equal to $1-s_{i\ell}$ and then to perform $k$-means clustering in this embedding. The resulting clustering fits a criterion similar to that of $RCut$.
@@
where $I$ is the identity matrix, and $\mathbf{1}$ is the (column) vector consisting of ones, both of appropriate dimensions, and $S$ is the already defined similarity matrix of $n$ objects. The matrix $A$ is non-negative and has a diagonal equal to zero, so it may be considered as a kind of (squared) pseudo-distance, needed by the Gower’s embedding method [Gower:1966]. Let $K$ be the $n\times n$ matrix of the (double centered) form [Gower:1966]:
@@
where $\Lambda$ is the diagonal matrix of the eigenvalues of $K$, and $V$ is the matrix whose columns are the corresponding eigenvectors (of unit length) of $K$. Denoting $Y=V\Lambda^{1/2}$, we obtain $K=YY^{T}$. According to Theorem 3, claim 4 in [MAK:2019:FI:Gower], detailing the respective theorem of [Gower:1966], the rows of $Y$, hence the columns of $Y^{T}$ constitute an embedding, called $K$-embedding, such that squared distances of embedded points are equal to squared distances, represented by the matrix AA. [^3] Note that the mentioned Theorem 3 was proven for Euclidean embeddings. For its validity under non-euclidean embeddings see [RAKMAKSTW:2020:trick].
@@
This differs slightly from the RCut criterion which seeks to minimize the average sum of similarities to elements outside of the cluster. The advantage is that it can be easily reformulated in terms of Term-Vector Space embedding so that cluster membership has a clear explanation in the textual domain, as shown in [Plosone2025].
@@ 2.2 $\mathcal{M}$-embedding
The $\mathcal{M}$-embedding, [Plosone2025], aims at placing objects $i,j$ at distances equal to $1/d_{ii}+1/d_{\ell\ell}-s_{i\ell}/(d_{ii}d_{\ell\ell})$, and then to perform weighted $k$-means clustering in this embedding. The resulting clustering fits a criterion identical to that of NCut. Let $\mathcal{E}$ be a matrix of the form
@@
Both NCut clustering and $\mathcal{M}$-embedding clustering seek to minimize the average sum of similarities to elements outside of the cluster, whereby the averaging is not performed via dividing by cluster cardinality but rather by cluster volume. The advantage is that $\mathcal{M}$-embedding can be easily reformulated in terms of term-vector space embedding so that cluster membership has a clear explanation in textual domain, as proven in [Plosone2025].
@@ 2.3 $\mathcal{B}$-embedding
This embedding can be easily reformulated in terms of term-vector space embedding so that cluster membership has a clear explanation in textual domain. For more details, see [RozdzialvMonografia:2025].
@@ 3 Coping with real datasets
The GSC methods perform very well when the similarity matrix fits the requirement of a block diagonal matrix, such that similarities of objects belonging to the same cluster (block) are on average higher than similarities between objects belonging to different clusters [Fischer:2005], [Su:2020]. Under such circumstances, for a broad range of similarity matrices, the clustering results using either method are (nearly) identical.
@@ 3.1 Review of literature on handling boundary data in clustering via rough sets
Rough set theory offers a range of tools to address uncertainty, vagueness, and incompleteness effectively. Among these tools, approximate clustering is an active area of research, see e.g. [Bello17] for a discussion and extensive bibliography. An interesting overview of applications of rough set theory in big data analysis and processing is given in [Pieta2021]. In particular, the authors discuss the advantages of combining this theory with other approaches, such as fuzzy sets, probabilistic concepts, and deep learning. They note that this latter hybrid idea seems to be promising for developing new methods, related tools, and extensions. See also [Peres:2012], especially the chapter by Lingras and Peters on pages 23–50, as well as Part XIV of the proceedings [Babu2014]. An easy and accessible introduction into various aspects of rough clustering is provided by  [Lingras2011]. Particularly these authors describe also methods hybridizing rough clustering with other techniques such as genetic algorithms, Kohonen self-organizing networks, and support vector clustering.
Since $k$-means clustering, is one of the most popular and effective clustering algorithms, see e.g. [Jain2010], we will focus on its different variants in the following.
@@
Here $X=(\mathbf{x}_{1},\dots,\mathbf{x}_{n})^{\mathrm{T}}$ is a set of $n$ objects described by real-valued vectors $\mathbf{x}_{i}$, $i=1,\dots,n$, $\boldsymbol{\mu}_{\ell}=(1/n_{\ell})\sum_{i\in C_{\ell}}\mathbf{x}_{i}$ is the centroid of $\ell$-th cluster, and $n_{\ell}$ is its size. The number $u_{i\ell}\in\{0,1\}$ equals 1 if $i$-th object belongs to cluster $C_{\ell}$ and $u_{i\ell}=0$ otherwise. $H\in\mathbb{R}^{n\times k}$ is a matrix with normalized degrees $h_{i\ell}=u_{i\ell}/\sqrt{n_{\ell}}$. Such a normalization ensures that $H$ is an orthonormal matrix, i.e. $H^{\texttt{T}}H=\mathbb{I}$. The first line of the above equation, (20a), explains the nature of HKM: it returns compact clusters containing elements that are as similar to each other as possible. A drawback of this algorithm is that it only returns satisfactory results if the clusters are linearly separable. The second line of the above equation, (20b), provides a solution. Noticing that $XX^{\texttt{T}}$ is a symmetric and semi-positive definite matrix (spd for short), we can replace $XX^{\texttt{T}}$ by any other spd matrix, $K$, which leads to a kernel $k$-mean, [STC04]. This new variant allows us to identify the non-linear structures, making it more suitable for real data sets. Formally, the kernel clustering algorithm relies on transforming samples from the original $m$-dimensional Euclidean space (where $m$ is the size of the vector $\mathbf{x}$) into a possibly infinite-dimensional feature space in which the differences between samples can be preserved and even enhanced. By using $k$-means for this new representation, we can get a better clustering performance. Of course, this requires choosing the right kernel function with the right parameters.
In practice, e.g. in text document clustering, we face situations where data can belong to several clusters simultaneously. To deal with such cases, [Bezdek81] generalized the objective function (20) by introducing the following criterion
@@
where $\tilde{u}_{i\ell}\in[0,1]$ is the degree to which element $\mathbf{x}_{i}$ belongs to cluster $C_{\ell}$, and $\alpha>1$ is a hyperparameter, called also fuzziness exponent. Typically, $\alpha=2$ is assumed. If $\alpha\rightarrow 1$, the values of $\tilde{u}_{i\ell}$ converge to 0 or 1, and the algorithm becomes identical to the $k$-means algorithm. On the other hand, if $\alpha\rightarrow\infty$, then the values of $\tilde{u}_{i\ell}\rightarrow 1/k$, i.e. we get a completely fuzzy partition. The algorithm minimizing the objective function (21) under additional probabilistic constraint $\sum_{\ell=1}^{k}\tilde{u}_{i\ell}=1$ is known as the Fuzzy $c$ Means, or FCM algorithm. Using membership values results in a more flexible and interpretable description of the clusters. In addition to its undoubted advantages, the algorithm also has significant drawbacks. Like KHM, it is sensitive to the choice of initial parameters and prefers spherical clusters. It also performs poorly on data sets with unbalanced clusters and is sensitive to noise and outliers. Furthermore, the computational complexity of FCM is higher than that of $k$-means, [axioms2024]. Various modifications that have been introduced to deal with these drawbacks are described, for example, in Section 3.3 of [STWMAKSpringer:2018].
@@
To broaden the applicability of FCM and improve the accuracy of clustering, two “kernelized” versions of FCM are developed: (a) centroids are constructed in the original $m$-dimensional Euclidean space, and (b) centroids are located in the (infinite-dimensional) feature space. Both approaches are briefly described by [GP10]. In addition, these authors conducted a comprehensive comparative analysis of fuzzy clustering and these two variants of kernel-based FCM. They found that there was no statistically significant difference between the kernel-based algorithms and FCM and its modification proposed by [Gustafson:1978]. They also observed that kernel-based FCM algorithms are generally amenable to the choice of specific values of the kernel parameters.
@@
FCM and HKM have another common feature: if the number of correctly grouped objects increases, the number of incorrectly grouped objects decreases. On the contrary, [Pet2015], [Peters2013] noted that in rough clustering, the numbers of correctly and incorrectly clustered objects are no longer complementary, hence the number of incorrectly clustered objects can be minimized independently. This property is important in practical applications, where minimizing the number of incorrectly clustered objects is more important than maximizing the number of correctly clustered objects.
@@
One of the first variants of rough $k$-means, called hereafter RKM, was proposed by [Lingras2004]. The authors assumed that each cluster $C_{\ell}$ is described by the lower, $\underline{C}_{\ell}$, and upper, $\overline{C}_{\ell}$, approximation. It is known, that $\underline{C}_{\ell}$ contains objects that certainly belong to $C_{\ell}$, while $\overline{C}_{\ell}\supseteq\underline{C}_{\ell}$ contains objects that possibly belong to $C_{\ell}$. Thus, the following properties must be satisfied:
@@
[Peters2006] has proposed some algorithm stability improvements. He modified the way of determining the sets $T(\mathbf{x})$ as well as the procedure for determining the cluster centroids. In this new formulation, these are computed as the convex combination of the centers of the lower and upper approximations, i.e.
@@
Another way to improve the quality of RKM by carefully choosing the hyperparameters $\tau$, $w^{L}$, and $w^{B}$ was proposed by [MITRA2004], who used a genetic algorithm to select appropriate values for these parameters. She assumed that $w^{L}\in(0.5,1.0)$ and $\tau\in(0,0.5)$. The Davies-Bouldin clustering validation index was used as the fitness function.
@@
[Zhang2016] modified approximate $k$-means by assigning different weights $w_{i\ell}$ to the features. Their solution is similar to the one used to compute centroids in FCM except that the weights are computed using a Gaussian function. The authors also describe related approaches to approximate clustering.
@@
Regarding the relationship between RKM and other variants of the $k$-means algorithm, [Peters2013] said that this algorithm is a useful enrichment of HKM and that it establishes a different subfamily (by processing uncertainty due to missing or incorrect information) compared to fuzzy and possibilistic clustering (which operate by processing ambiguity based on similarity). This is in agreement with the known fact that fuzzy sets and rough sets are orthogonal concepts, [Pawlak1985]. However, [Dubois1990] proposed a different perspective on this problem. Accepting the fact that the two tools are designed for different tasks, they stated that their joint use can bring many advantages. This has given rise to new directions referred to as “fuzzy rough” and “rough fuzzy”. In the case of clustering, we should mention the paper [Mitra2006], which proposed an extension of RKM to the rough fuzzy $c$-means, or RFCM for short. Here, each sample $\mathbf{x}_{i}$ occurring in the equation (22) is weighted by the fuzzy membership value  $\tilde{u}_{i\ell}^{\alpha}$ ($\alpha=2$), similarly to the FCM algorithm. An additional value of this paper is that the authors described in detail how to calculate the rough and rough fuzzy variants of the Davies-Bouldin and Dunn indices used to assess clustering quality. The main focus of this article was the adaptation of a new clustering architecture, referred to as collaborative clustering and introduced by [Pedrycz2002]. Briefly, the method involves processing several subsets of patterns together to find a common structure for all of them, see, for example,  [Cornuejols2018] for a deeper explanation.
@@
Another collaborative approaches were proposed by [HU:2017, Yue:2023]. [HU:2017] combine ensemble clustering with supportive random forest classifiers. [Yue:2023] concentrate on cluster intersection of clusterings obtained using different base clustering methods. They define the similarity of objects as the counts of their co-occurrence in clusters in base clustering (co-association matrix consists of these counts). They then apply the standard Graph Spectral Clustering method to the co-association matrix.
@@
[Maji2007] noted a weakness of the RFCM: each object has a distinct weight (its fuzzy membership value), resulting in shifting centroids from their desired locations. The algorithm is also sensitive to noise and outliers. The authors proposed the RFPCM (rough-fuzzy possibilistic $c$-means) algorithm to improve it. They postulate that all objects in a lower approximation should have an identical influence on the determination of their centers. This way they developed an algorithm with crisp lower approximations and fuzzy boundaries. Later, [Paul2014], proposed so-called robust rough-fuzzy $c$-means, in which each cluster is described by its centroid (prototype), a possibilistic fuzzy lower approximation, and a probabilistic fuzzy (upper) boundary. The centroid depends on the mentioned lower approximation and the upper boundary. The algorithm detects overlapping clusters with arbitrary shapes in the presence of noise.
@@
Recently, [Zhao2023] proposed a more sophisticated approach to this problem. First, the so-called transfer learning mechanism, see [Zhuang2021], is introduced into rough clustering. Then the transfer rough clustering algorithm’s objective function is optimized using the differential evolution algorithm.
@@
[Manish2010] carried out a comparative study of fuzzy and rough clustering and found out that though a conversion between the results of both is frequently possible, the rough set-based clustering has the advantage of better descriptiveness in high dimensions. The advantages of the rough set approach are exploited to improve the FCM algorithm by [YU2024], leading to creation of the FRCM algorithm (fuzzy-rough $c$-means).

### v2

# arXiv:2512.12436v2[cs.LG] 27 Feb 2026

@@ 1 Introduction
Explainability turns out to be crucial for the application of Machine Learning techniques, in particular clustering. From this perspective, Graph Spectral Clustering (GSC for short) methods have some serious drawbacks as they operate in a space hard to understand by humans, that is, in the Laplacian spectral domain, and this is why they are considered opaque models [^1] See also, for example [41], and https://crunchingthedata.com/when-to-use-spectral-clustering/. . Model-unaware methods of textual clustering are known and could be considered an option, but they only describe groups and do not explain why documents are included. Recently explanatory clustering methods for Combinatorial and Normalized Laplacian-based GSC were proposed for the textual domain, as presented in [45]. Yet, two problems remain open: (1) in real-world settings, there exist objects that do not fit any cluster and are forced into some clusters just distorting them, and (2) clustering methods do not achieve the real optimal value of the quality function so that some cluster elements may also not fit the concepts and can distort them. In this paper, we want to contribute to solving both problems by applying concepts from rough set theory, where entities, e.g. documents, are divided into core and boundary objects.
@@
Thus, the approach introduced in the context of rough sets by [31] may prove fruitful here as it allows one to distinguish between cluster core elements and boundary elements. This distinction can, for example, result from repeated runs of a particular probabilistic clustering algorithm. Elements that belong to the same cluster on each run are (with high probability) the core of a cluster, while the others can be considered to belong to the boundaries of different clusters. Another situation, though bearing some similarity, is when we apply different algorithms optimizing the same criterion, but each in an approximate way. The elements fitting together may be considered the core, with a high probability of fitting the (approximated) criterion, while the other will constitute boundaries of appropriate clusters. However, another approach can be a robustness study, checking the impact on the optimality criterion of randomly shuffling parts of the established clusters.
@@ 2 Brief introduction to Graph Spectral Analysis
There are several graph clustering methods; see, for example, [42], aimed at discovering clusters that meet related but slightly different criteria. The best known are the RCut criterion and the NCut criterion.
Assume that the matrix $S$, called the similarity matrix, represents similarities between pairs of $n$ objects with entries ranging from 0 to 1. $S$ is assumed to have a diagonal equal to zero, as self-loops are not allowed (see, e.g. [49]).
@@
where $n_{j}=|C_{j}|$. A detailed discussion leading to the above formulation can be found, for example, in [45].
@@
where $\mathcal{V}_{j}=\sum_{i\in C_{j}}d_{ii}$ denotes the volume of the node $j$, and $d_{ii}=\sum_{j=1}^{n}s_{ij}$ stands for the degree of the node $i$. Thus, we try to minimize the total weights of links joining nodes belonging to opposite clusters. In the first case, this sum is divided by the cardinalities of the respective clusters, while in the second case - by the volumes of these clusters. Again, see [45] for a more in-depth discussion.
@@
where $D$ is the degree matrix, i.e., a diagonal matrix whose elements are the degrees $d_{jj}$ of all nodes. A normalized Laplacian [^2] Other Laplacians are also used (see, [54]), e.g. the random walk Laplacian $\mathbb{L}$ of a graph, defined as $\mathbb{L}=LD^{-1}=I-SD^{-1}$ (5) $\mathcal{L}$ of the graph represented by $S$ is defined as
@@
Based on the Rayleigh-Ritz theorem [11], it is concluded that the sought pseudo-indicator vectors are the eigenvectors of the respective Laplacian, related to the lowest eigenvalues. If we cluster into $k$ clusters, we take the lowest $k$ eigenvectors (except the last, which is certain to be equal to zero). These pseudo-indicator vectors span the embedding space for the objects which are then clustered in this embedding using, e.g. $k$-means algorithm. For an explanation and further details, see, e.g. [51] or [54]. Afterward, note that $L$-based clustering means GSC based on combinatorial Laplacian $L$, and $N$-based clustering means GSC based on normalized Laplacian $\mathfrak{L}$. Note that by projecting the original data onto the space spanned by $k$ eigenvectors of the corresponding Laplacian, we discover an interesting feature of such a procedure: it naturally combines clustering with (nonlinear) dimensionality reduction.
@@ 2.1 $K$-embedding
The $K$-embedding [45] aims at placing objects $i,\ell$ at distances equal to $1-s_{i\ell}$ and then to perform $k$-means clustering in this embedding. The resulting clustering fits a criterion similar to that of $RCut$.
@@
where $I$ is the identity matrix, and $\mathbf{1}$ is the (column) vector consisting of ones, both of appropriate dimensions, and $S$ is the already defined similarity matrix of $n$ objects. The matrix $A$ is non-negative and has a diagonal equal to zero, so it may be considered as a kind of (squared) pseudo-distance, needed by the Gower’s embedding method [12]. Let $K$ be the $n\times n$ matrix of the (double centered) form [12]:
@@
where $\Lambda$ is the diagonal matrix of the eigenvalues of $K$, and $V$ is the matrix whose columns are the corresponding eigenvectors (of unit length) of $K$. Denoting $Y=V\Lambda^{1/2}$, we obtain $K=YY^{T}$. According to Theorem 3, claim 4 in [20], detailing the respective theorem of [12], the rows of $Y$, hence the columns of $Y^{T}$ constitute an embedding, called $K$-embedding, such that squared distances of embedded points are equal to squared distances, represented by the matrix AA. [^3] Note that the mentioned Theorem 3 was proven for Euclidean embeddings. For its validity under non-euclidean embeddings see [21].
@@
This differs slightly from the RCut criterion which seeks to minimize the average sum of similarities to elements outside of the cluster. The advantage is that it can be easily reformulated in terms of Term-Vector Space embedding so that cluster membership has a clear explanation in the textual domain, as shown in [45].
@@ 2.2 $\mathcal{M}$-embedding
The $\mathcal{M}$-embedding, [45], aims at placing objects $i,j$ at distances equal to $1/d_{ii}+1/d_{\ell\ell}-s_{i\ell}/(d_{ii}d_{\ell\ell})$, and then to perform weighted $k$-means clustering in this embedding. The resulting clustering fits a criterion identical to that of NCut. Let $\mathcal{E}$ be a matrix of the form
@@
Both NCut clustering and $\mathcal{M}$-embedding clustering seek to minimize the average sum of similarities to elements outside of the cluster, whereby the averaging is not performed via dividing by cluster cardinality but rather by cluster volume. The advantage is that $\mathcal{M}$-embedding can be easily reformulated in terms of term-vector space embedding so that cluster membership has a clear explanation in textual domain, as proven in [45].
@@ 2.3 $\mathcal{B}$-embedding
This embedding can be easily reformulated in terms of term-vector space embedding so that cluster membership has a clear explanation in textual domain. For more details, see [46].
@@ 3 Coping with real datasets
The GSC methods perform very well when the similarity matrix fits the requirement of a block diagonal matrix, such that similarities of objects belonging to the same cluster (block) are on average higher than similarities between objects belonging to different clusters [10], [47]. Under such circumstances, for a broad range of similarity matrices, the clustering results using either method are (nearly) identical.
@@ 3.1 Review of literature on handling boundary data in clustering via rough sets
Rough set theory offers a range of tools to address uncertainty, vagueness, and incompleteness effectively. Among these tools, approximate clustering is an active area of research, see e.g. [3] for a discussion and extensive bibliography. An interesting overview of applications of rough set theory in big data analysis and processing is given in [40]. In particular, the authors discuss the advantages of combining this theory with other approaches, such as fuzzy sets, probabilistic concepts, and deep learning. They note that this latter hybrid idea seems to be promising for developing new methods, related tools, and extensions. See also [39], especially the chapter by Lingras and Peters on pages 23–50, as well as Part XIV of the proceedings [2]. An easy and accessible introduction into various aspects of rough clustering is provided by  [23]. Particularly these authors describe also methods hybridizing rough clustering with other techniques such as genetic algorithms, Kohonen self-organizing networks, and support vector clustering.
Since $k$-means clustering, is one of the most popular and effective clustering algorithms, see e.g. [18], we will focus on its different variants in the following.
@@
Here $X=(\mathbf{x}_{1},\dots,\mathbf{x}_{n})^{\mathrm{T}}$ is a set of $n$ objects described by real-valued vectors $\mathbf{x}_{i}$, $i=1,\dots,n$, $\boldsymbol{\mu}_{\ell}=(1/n_{\ell})\sum_{i\in C_{\ell}}\mathbf{x}_{i}$ is the centroid of $\ell$-th cluster, and $n_{\ell}$ is its size. The number $u_{i\ell}\in\{0,1\}$ equals 1 if $i$-th object belongs to cluster $C_{\ell}$ and $u_{i\ell}=0$ otherwise. $H\in\mathbb{R}^{n\times k}$ is a matrix with normalized degrees $h_{i\ell}=u_{i\ell}/\sqrt{n_{\ell}}$. Such a normalization ensures that $H$ is an orthonormal matrix, i.e. $H^{\texttt{T}}H=\mathbb{I}$. The first line of the above equation, (20a), explains the nature of HKM: it returns compact clusters containing elements that are as similar to each other as possible. A drawback of this algorithm is that it only returns satisfactory results if the clusters are linearly separable. The second line of the above equation, (20b), provides a solution. Noticing that $XX^{\texttt{T}}$ is a symmetric and semi-positive definite matrix (spd for short), we can replace $XX^{\texttt{T}}$ by any other spd matrix, $K$, which leads to a kernel $k$-mean, [43]. This new variant allows us to identify the non-linear structures, making it more suitable for real data sets. Formally, the kernel clustering algorithm relies on transforming samples from the original $m$-dimensional Euclidean space (where $m$ is the size of the vector $\mathbf{x}$) into a possibly infinite-dimensional feature space in which the differences between samples can be preserved and even enhanced. By using $k$-means for this new representation, we can get a better clustering performance. Of course, this requires choosing the right kernel function with the right parameters.
In practice, e.g. in text document clustering, we face situations where data can belong to several clusters simultaneously. To deal with such cases, [4] generalized the objective function (20) by introducing the following criterion
@@
where $\tilde{u}_{i\ell}\in[0,1]$ is the degree to which element $\mathbf{x}_{i}$ belongs to cluster $C_{\ell}$, and $\alpha>1$ is a hyperparameter, called also fuzziness exponent. Typically, $\alpha=2$ is assumed. If $\alpha\rightarrow 1$, the values of $\tilde{u}_{i\ell}$ converge to 0 or 1, and the algorithm becomes identical to the $k$-means algorithm. On the other hand, if $\alpha\rightarrow\infty$, then the values of $\tilde{u}_{i\ell}\rightarrow 1/k$, i.e. we get a completely fuzzy partition. The algorithm minimizing the objective function (21) under additional probabilistic constraint $\sum_{\ell=1}^{k}\tilde{u}_{i\ell}=1$ is known as the Fuzzy $c$ Means, or FCM algorithm. Using membership values results in a more flexible and interpretable description of the clusters. In addition to its undoubted advantages, the algorithm also has significant drawbacks. Like KHM, it is sensitive to the choice of initial parameters and prefers spherical clusters. It also performs poorly on data sets with unbalanced clusters and is sensitive to noise and outliers. Furthermore, the computational complexity of FCM is higher than that of $k$-means, [35]. Various modifications that have been introduced to deal with these drawbacks are described, for example, in Section 3.3 of [54].
@@
To broaden the applicability of FCM and improve the accuracy of clustering, two “kernelized” versions of FCM are developed: (a) centroids are constructed in the original $m$-dimensional Euclidean space, and (b) centroids are located in the (infinite-dimensional) feature space. Both approaches are briefly described by [13]. In addition, these authors conducted a comprehensive comparative analysis of fuzzy clustering and these two variants of kernel-based FCM. They found that there was no statistically significant difference between the kernel-based algorithms and FCM and its modification proposed by [15]. They also observed that kernel-based FCM algorithms are generally amenable to the choice of specific values of the kernel parameters.
@@
FCM and HKM have another common feature: if the number of correctly grouped objects increases, the number of incorrectly grouped objects decreases. On the contrary, [37], [38] noted that in rough clustering, the numbers of correctly and incorrectly clustered objects are no longer complementary, hence the number of incorrectly clustered objects can be minimized independently. This property is important in practical applications, where minimizing the number of incorrectly clustered objects is more important than maximizing the number of correctly clustered objects.
@@
One of the first variants of rough $k$-means, called hereafter RKM, was proposed by [22]. The authors assumed that each cluster $C_{\ell}$ is described by the lower, $\underline{C}_{\ell}$, and upper, $\overline{C}_{\ell}$, approximation. It is known, that $\underline{C}_{\ell}$ contains objects that certainly belong to $C_{\ell}$, while $\overline{C}_{\ell}\supseteq\underline{C}_{\ell}$ contains objects that possibly belong to $C_{\ell}$. Thus, the following properties must be satisfied:
@@
[36] has proposed some algorithm stability improvements. He modified the way of determining the sets $T(\mathbf{x})$ as well as the procedure for determining the cluster centroids. In this new formulation, these are computed as the convex combination of the centers of the lower and upper approximations, i.e.
@@
Another way to improve the quality of RKM by carefully choosing the hyperparameters $\tau$, $w^{L}$, and $w^{B}$ was proposed by [28], who used a genetic algorithm to select appropriate values for these parameters. She assumed that $w^{L}\in(0.5,1.0)$ and $\tau\in(0,0.5)$. The Davies-Bouldin clustering validation index was used as the fitness function.
@@
[57] modified approximate $k$-means by assigning different weights $w_{i\ell}$ to the features. Their solution is similar to the one used to compute centroids in FCM except that the weights are computed using a Gaussian function. The authors also describe related approaches to approximate clustering.
@@
Regarding the relationship between RKM and other variants of the $k$-means algorithm, [38] said that this algorithm is a useful enrichment of HKM and that it establishes a different subfamily (by processing uncertainty due to missing or incorrect information) compared to fuzzy and possibilistic clustering (which operate by processing ambiguity based on similarity). This is in agreement with the known fact that fuzzy sets and rough sets are orthogonal concepts, [32]. However, [9] proposed a different perspective on this problem. Accepting the fact that the two tools are designed for different tasks, they stated that their joint use can bring many advantages. This has given rise to new directions referred to as “fuzzy rough” and “rough fuzzy”. In the case of clustering, we should mention the paper [27], which proposed an extension of RKM to the rough fuzzy $c$-means, or RFCM for short. Here, each sample $\mathbf{x}_{i}$ occurring in the equation (22) is weighted by the fuzzy membership value  $\tilde{u}_{i\ell}^{\alpha}$ ($\alpha=2$), similarly to the FCM algorithm. An additional value of this paper is that the authors described in detail how to calculate the rough and rough fuzzy variants of the Davies-Bouldin and Dunn indices used to assess clustering quality. The main focus of this article was the adaptation of a new clustering architecture, referred to as collaborative clustering and introduced by [33]. Briefly, the method involves processing several subsets of patterns together to find a common structure for all of them, see, for example,  [7] for a deeper explanation.
@@
Another collaborative approaches were proposed by [16, 56]. [16] combine ensemble clustering with supportive random forest classifiers. [56] concentrate on cluster intersection of clusterings obtained using different base clustering methods. They define the similarity of objects as the counts of their co-occurrence in clusters in base clustering (co-association matrix consists of these counts). They then apply the standard Graph Spectral Clustering method to the co-association matrix.
@@
[25] noted a weakness of the RFCM: each object has a distinct weight (its fuzzy membership value), resulting in shifting centroids from their desired locations. The algorithm is also sensitive to noise and outliers. The authors proposed the RFPCM (rough-fuzzy possibilistic $c$-means) algorithm to improve it. They postulate that all objects in a lower approximation should have an identical influence on the determination of their centers. This way they developed an algorithm with crisp lower approximations and fuzzy boundaries. Later, [30], proposed so-called robust rough-fuzzy $c$-means, in which each cluster is described by its centroid (prototype), a possibilistic fuzzy lower approximation, and a probabilistic fuzzy (upper) boundary. The centroid depends on the mentioned lower approximation and the upper boundary. The algorithm detects overlapping clusters with arbitrary shapes in the presence of noise.
@@
Recently, [58] proposed a more sophisticated approach to this problem. First, the so-called transfer learning mechanism, see [60], is introduced into rough clustering. Then the transfer rough clustering algorithm’s objective function is optimized using the differential evolution algorithm.
@@
[19] carried out a comparative study of fuzzy and rough clustering and found out that though a conversion between the results of both is frequently possible, the rough set-based clustering has the advantage of better descriptiveness in high dimensions. The advantages of the rough set approach are exploited to improve the FCM algorithm by [55], leading to creation of the FRCM algorithm (fuzzy-rough $c$-means).